Fermi National Laboratory


Accelerator Update - July 23-July 25

The integrated luminosity for last week was 5106.9 nb-1.

Wednesday July 23
The day shift began with Operations monitoring store 2813, with the Antiproton Source (Pbar) stacking, with Switchyard and Recycler experts conducting beam-commissioning studies, and with MiniBooNE taking beam.

Operators and Switchyard experts established beam to the Switchyard dump at 8:10 AM.

The Linac 400 MeV area suffered six trips between 8:17 AM and 8:42 AM. Experts investigated.

A Main Injector expert began a slipstacking study at 9:19 AM.

A Booster RF station (BRF6) tripped off due to a high vacuum indication at 12:41 PM.

MiniBooNE suffered the first of three beam trips in ten minutes at 2:55 PM. Then it tripped off two more times at 3:07 PM.

At 3:21 PM, the Tevatron (TeV) suffered a sector D4 quench. A cryo system control crate power supply tripped off due to a blown fuse. An expert replaced the power supply.

Switchyard and Recycler experts completed their studies at 5:25 PM.

At 5:56 PM, a Booster power supply (VCB1) tripped off due to an over-temperature problem. It was back on nine minutes later.

Operations put the TeV into shot setup at 7:46 PM, and at 10:32 PM established store 2815 with an initial luminosity of 38.63E30. Pbar resumed stacking.

Thursday July 24
The midnight shift began with Operations monitoring store 2815, with Pbar stacking, and with MiniBooNE taking beam.

At 2:30 AM, MiniBooNE asked to have their beam turned off while they fixed a Data Acquisition problem. It wasn't until 6:55 AM that they resumed taking beam.

The Booster VCB1 power supply tripped off at 3:43 AM and then again at 4:53 AM. It was reset within a few minutes after each trip.

Two Booster RF stations (BRF5 & 6) tripped off at 6:52 AM due to a high vacuum indication.

Operations turned off beam at 8 AM to MiniBooNE for a scheduled access into MI12 and 13.

A TeV sector A0 cryo compressor tripped off at 9:30 AM. Cryo technicians investigated.

MiniBooNE resumed taking beam at 12:25 PM.

At 1:09 PM, a number of vacuum pumps tripped off and vacuum valves closed in the A1 transfer line due to a vacuum burst. A half hour later an expert reported that the vacuum was looking better. He thought that the pumps had overheated and started outgassing. The valves were still closed. They will be opened when the store ends.

At 4:07 PM, a water heater fuse blew and tripped off a TeV RF station (TRF8). A few minutes later TRF5, 6, & 7 all tripped off due to vacuum burst. This caused store 2815 to abort.

Operator prepared D-zero (D0) for access at 5:17 PM. They came out of access at 6 PM.

Operations put the TeV through a dry squeeze at 7:06 PM.

Operations put the TeV into shot setup at 7:35 PM.

At 7:58 PM, A TeV RF station (TRF4) tripped off. Operators investigated and consulted with an expert.

The Booster VCB1 power supply tripped again at 8:29 PM. It reset a few minutes later.

Operation established store 2817 at 10:34 PM with an initial luminosity of 42.64E30. Pbar resumed stacking. (The highest luminosity set this run was back on 5/17/03 with a luminosity of 44.89E30.)

Friday July 25
The midnight shift began with Operations monitoring store 2817, with Pbar stacking, and with MiniBooNE taking beam.

The Booster VCB1 power supply tripped again at 1:20 AM. It reset a few minutes later.

At 6:42 AM, the I- Source suffered a short period of sparking. Operators lowered its high voltage very slightly, which seemed to help.

Today's Plan
The plan for today is to put the TeV into shot setup around 7 PM tonight. A credible source has said that the Fermilab accelerator complex will go into a ten-week shutdown starting September 8th.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/25/2003   email Fermilab